# Value-Incentivized Preference Optimization: A Unified Approach to Online and Offline RLHF

### v1

@@ 1.2 Related work
The principles of optimism and pessimism are typically implemented via constructing confidence intervals or posterior sampling, which have been demonstrated to be provably efficient in tabular settings (Jin et al.,, 2018; Shi et al.,, 2022). Yet, these approaches have had limited success in conjunction with deep learning architectures (Gawlikowski et al.,, 2021), and many empirical heuristics in turn lack theoretical validation (Kumar et al.,, 2020). VPO draws inspiration from reward-biased exploration (Kumar and Becker,, 1982; Liu et al.,, 2020; Hung et al.,, 2021; Mete et al.,, 2021) in the standard online RL literature, but significantly broadens its scope to the offline setting and RLHF for the first time.
@@ 2 Preliminaries
where $y_{1}\succ y_{2}$ indicates that $y_{1}$ is preferred over $y_{2}$, $r^{\star}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ is the ground truth reward function, and $\sigma:\mathbb{R}\to(0,1)$ is the logistic function. A preference data sample is denoted by a tuple $(x,y_{+},y_{-})$, where $y_{+}$ (resp. $y_{0}$) is the preferred (resp. unpreferred) answer in the comparison.
@@ 3.1 General framework
In view of the sub-optimality of naive MLE for reward estimation (Cesa-Bianchi et al.,, 2017; Rashidinejad et al.,, 2022), and motivated by the effectiveness of reward-biased MLE in online RL (Kumar and Becker,, 1982; Liu et al.,, 2020), we propose to regularize the reward estimate via
@@
At first glance, the objective function for VPO (12) does not immediately imply a computationally-efficient algorithm due to the presence of $J^{\star}(r)$. However, by exploiting the same closed-form solution for the optimal policy given the reward in (6), and the reward representation inferred from the policy vai (8), we can explicitly express $J^{\star}(r)$ as
@@ 3.2 Online RLHF: algorithm and theory
The next theorem demonstrates that Algorithm 1 achieves $\widetilde{\mathcal{O}}(\sqrt{T})$ cumulative regret under mild assumptions. The proof is provided in Appendix A.
@@ 3.3 Offline RLHF: algorithm and theory
Under Assumptions 1 and 2, let $r_{\widehat{\theta}}\in\Theta$ denote the corresponding reward model for $\widehat{\pi}$. Assume that $\|\theta^{\star}\|_{2}\leq C$ and $\|\widehat{\theta}\|_{2}\leq C$ for some $C>0$. Let $\alpha=\sqrt{N}$ and $\delta\in(0,1)$. With probability $1-\delta$, we have
@@ A.2 Proof of Lemma 2
Lemma 5 ((Zhang,, 2023, Theorem 13.2))

### v2

@@ 1.2 Related work
The principles of optimism and pessimism are typically implemented via constructing confidence intervals or posterior sampling, which have been demonstrated to be provably efficient in tabular settings (Jin et al.,, 2018; Shi et al.,, 2022). Yet, these approaches have had limited success in conjunction with deep learning architectures (Gawlikowski et al.,, 2021), and many empirical heuristics in turn lack theoretical validation (Kumar et al.,, 2020). VPO draws inspiration from reward-biased exploration (Kumar and Becker,, 1982; Liu et al.,, 2020, 2024; Hung et al.,, 2021; Mete et al.,, 2021) in the standard online RL literature, but significantly broadens its scope to the offline setting and RLHF for the first time.
@@ 2 Preliminaries
where $y_{1}\succ y_{2}$ indicates that $y_{1}$ is preferred over $y_{2}$, $r^{\star}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ is the ground truth reward function, and $\sigma:\mathbb{R}\to(0,1)$ is the logistic function. A preference data sample is denoted by a tuple $(x,y_{+},y_{-})$, where $y_{+}$ (resp. $y_{-}$) is the preferred (resp. unpreferred) answer in the comparison.
@@ 3.1 General framework
In view of the sub-optimality of naive MLE for reward estimation (Cesa-Bianchi et al.,, 2017; Rashidinejad et al.,, 2022), and motivated by the effectiveness of reward-biased MLE in online RL (Kumar and Becker,, 1982; Liu et al.,, 2020, 2024), we propose to regularize the reward estimate via
@@
At first glance, the objective function for VPO (12) does not immediately imply a computationally-efficient algorithm due to the presence of $J^{\star}(r)$. However, by exploiting the same closed-form solution for the optimal policy given the reward in (6), and the reward representation inferred from the policy via (8), we can explicitly express $J^{\star}(r)$ as
@@ 3.2 Online RLHF: algorithm and theory
The next theorem demonstrates that Algorithm 1 achieves $\widetilde{\mathcal{O}}(\sqrt{T})$ cumulative regret under mild assumptions. The proof is provided in Appendix A. The proof logic follows from that of (Liu et al.,, 2024).
@@ 3.3 Offline RLHF: algorithm and theory
Under Assumptions 1 and 2, let $\widehat{\theta}\in\Theta$ denote the corresponding reward model for $\widehat{\pi}$. Assume that $\|\theta^{\star}\|_{2}\leq C$ and $\|\widehat{\theta}\|_{2}\leq C$ for some $C>0$. Let $\alpha=\sqrt{N}$ and $\delta\in(0,1)$. With probability $1-\delta$, we have
@@
4 Token-level VPO
Recently, Rafailov et al., (2024) offered an interpretation of DPO using the token-level Markov Decision Process (MDP), aiming at reconciling the gap between the practical fine-tuning of LLMs at the token level and the theoretical formulation of DPO at the sentence level. Fortunately, this interpretation does not require algorithmic modifications. Below we provide a short discussion and highlight that VPO can be similarly interpreted using the token-level MDP.
Token-level MDP and preference modeling.
Recall that in LLMs, the prompt $x$ can be broken into a sequence of tokens, e.g., $x=(x_{0},\ldots,x_{m})$, from a fixed discrete vocabulary $\mathcal{A}$. We define the token-level MDP as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r^{\star},H)$, where $H$ is the horizon length, i.e., the longest possible number of tokens in a sentence. The state space $\mathcal{S}$ consists of all the possible token combinations of length $H$, and the transition kernel is deterministically defined as follows.
  1. 1.
    The initial state is defined by the prompt $x$ as $s_{0}=\{x_{0},\cdots,x_{m}\}$;
  2. 2.
    Given the response $y=\{y_{0},\cdots,y_{i-1}\}$ up to the $i$-th token, the state at step $i$ is defined as $s_{i}=\{x_{0},\cdots,x_{m},y_{0},\cdots,y_{i-1}\}$;
  3. 3.
    Upon an action of the LLM for the next token $a_{i}=y_{i}$, the next state at the token-level MDP deterministically becomes $s_{i+1}=(s_{i},a_{i})=(x_{0},\cdots,x_{m},y_{0},\cdots,y_{i})$.
We assume that the last token of a sentence, the EOS token, is absorbing, such that the token-level MDP stays in the corresponding state as soon as the last action is the EOS token. With slight abuse of notation from earlier sections, the reward function $r^{\star}(s,a)$ defines the ground truth reward at state $s$ upon action $a$.
Given a pair of trajectories $\tau_{1}=\{s_{0},a_{0}^{1},\ldots,s_{H-1}^{1},a_{H-1}^{1},s_{H}^{1}\}$ and $\tau_{2}=\{s_{0},a_{0}^{2},\ldots,s_{H-1}^{2},a_{H-1}^{2},s_{H}^{2}\}$, the corresponding Bradley-Terry preference model (Bradley and Terry,, 1952) is
$\mathbb{P}(\tau_{1}\succ\tau_{2})$ $=\frac{\exp\left(\sum_{i=0}^{H-1}r^{\star}(s_{i}^{1},a_{i}^{1})\right)}{\exp\left(\sum_{i=0}^{H-1}r^{\star}(s_{i}^{1},a_{i}^{1})\right)+\exp\left(\sum_{i=0}^{H-1}r^{\star}(s_{i}^{2},a_{i}^{2})\right)}$
$=\sigma\left(\sum_{i=0}^{H-1}r^{\star}(s_{i}^{1},a_{i}^{1})-\sum_{i=0}^{H-1}r^{\star}(s_{i}^{2},a_{i}^{2})\right),$
A preference data sample is denoted by a tuple $(x,\tau_{+},\tau_{-})$, where $\tau_{+}$ (resp. $\tau_{-}$) is the preferred (resp. unpreferred) answer in the comparison. Given a preference dataset $\mathcal{D}$ composed of independent samples, The negative log-likelihood can be defined as
$\ell(r,\mathcal{D})\coloneqq-\sum_{(\tau_{+},\tau_{-})\in\mathcal{D}}\log\sigma\left(\sum_{i=0}^{H-1}r(s_{i}^{+},a_{i}^{+})-\sum_{i=0}^{H-1}r(s_{i}^{-},a_{i}^{-})\right),$ (25)
where $(s_{i}^{+},a_{i}^{+})$ (resp. $(s_{i}^{-},a_{i}^{-})$) are the state-action pairs in the trajectory $\tau_{+}$ (resp. $\tau_{-}$).
Token-level RL fine-tuning.
Let the entropy of policy $\pi$ under initial state distribution $s_{0}\sim\rho$ be defined as
$\mathcal{H}(\rho,\pi)\coloneqq-\underset{s_{0}\sim\rho,\atop a_{i}\sim\pi(\cdot|s_{i})}{{\mathbb{E}}}\left[\sum_{i=0}^{H-1}\log\pi(a_{i}|s_{i})\right],$
and $\mathcal{H}(s,\pi)$ be the entropy when the initial state $s_{0}=s$. Given a reward function $r$, we define the KL-constrained RL objective against the reference policy $\pi_{\text{ref}}$ as (Rafailov et al.,, 2024):
$\pi_{r}\coloneqq\mathop{\mathrm{argmax}}_{\pi}J(r,\pi)$ $\coloneqq\underset{s_{0}\sim\rho,\atop a_{i}\sim\pi(\cdot|s_{i})}{{\mathbb{E}}}\Big{[}\sum_{i=0}^{H-1}(\underbrace{r(s_{i},a_{i})+\beta\log\pi_{\text{ref}}(a_{i}|s_{i})}_{r_{\beta}(s_{i},a_{i})})\Big{]}+\beta\mathcal{H}(\rho,\pi),$ (26)
where we denote $r_{\beta}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ as follows:
$\forall(s,a)\in\mathcal{S}\times\mathcal{A}:\quad r_{\beta}(s,a)\coloneqq r(s,a)+\beta\log\pi_{\text{ref}}(a|s),$ (27)
which can be seen as the actual token-wise reward function optimized by the LLM.
Token-level DPO.
The KL-constrained RL objective (26) has a closed-form solution (Nachum et al.,, 2017; Cen et al.,, 2022) given by
$\forall(s,a)\in\mathcal{S}\times\mathcal{A}:\quad\pi_{r}(a|s)=\exp((Q_{\beta}^{\star}(s,a)-V_{\beta}^{\star}(s))/\beta),$ (28)
where $V_{\beta}^{\star}:\mathcal{S}\rightarrow\mathbb{R}$ and $Q_{\beta}^{\star}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ are the optimal soft value and Q functions, respectively,
$\forall s\in\mathcal{S}:\quad V_{\beta}^{\star}(s)\coloneqq\underset{a_{i}\sim\pi_{r}(\cdot|s_{i})}{{\mathbb{E}}}\left[\sum_{i=0}^{H-1}r_{\beta}(s_{i},a_{i})|s_{0}=s\right]+\beta\mathcal{H}(s,\pi_{r})$ (29)
denote the optimal soft value function w.r.t. the reward function $r_{\beta}$, and
$\forall(s,a)\in\mathcal{S}\times\mathcal{A}:\quad Q_{\beta}^{\star}(s,a)\coloneqq r_{\beta}(s,a)+V_{\beta}^{\star}(s^{\prime}),$ (30)
where $s^{\prime}=(s,a)$ is the deterministic next state. Plugging (30) and (27) into (28) implies that, for any trajectory $\tau=\{s_{0},a_{0},\cdots,a_{H-1},s_{H}\}$, Rafailov et al., (2024) shows
$\sum_{i=0}^{H-1}r(s_{i},a_{i})$ $=\sum_{i=0}^{H-1}\left(Q_{\beta}^{\star}(s_{i},a_{i})-\beta\log\pi_{\text{ref}}(a_{i}|s_{i})-V_{\beta}^{\star}(s_{i+1})\right)$
$=Q_{\beta}^{\star}(s_{0},a_{0})-\beta\log\pi_{\text{ref}}(a_{0}|s_{0})+\sum_{i=1}^{H-1}\left(Q_{\beta}^{\star}(s_{i},a_{i})-V_{\beta}^{\star}(s_{i})-\beta\log\pi_{\text{ref}}(a_{i}|s_{i})\right)$
$=V_{\beta}^{\star}(s_{0})+\beta\sum_{i=0}^{H-1}\log\frac{\pi_{r}(a_{i}|s_{i})}{\pi_{\text{ref}}(a_{i}|s_{i})},$ (31)
where the second line uses $V_{\beta}^{\star}(s_{H})=0$ at the terminal state. Thus the DPO loss (which is the negative log-likelihood loss) could be written as
$\mathcal{L}(\pi,\mathcal{D})=-\sum_{(\tau_{+},\tau_{-})\in\mathcal{D}}\log\sigma\left(\beta\sum_{i=0}^{H-1}\log\frac{\pi(a_{i}^{+}|s_{i}^{+})}{\pi_{\text{ref}}(a_{i}^{+}|s_{i}^{+})}-\beta\sum_{i=0}^{H-1}\log\frac{\pi(a_{i}^{-}|s_{i}^{-})}{\pi_{\text{ref}}(a_{i}^{-}|s_{i}^{-})}\right).$ (32)
Token-level VPO.
With slight abuse of notation, define
$J^{\star}(r)\coloneqq\max_{\pi}J(r,\pi),$ (33)
which is used as the bias term in regularizing the reward estimation in VPO. Again, we impose the following assumption to deal with the shift ambiguity issue caused by the Bradley-Terry model:
Assumption 3
We assume that $r^{\star}\in\mathcal{R}$, where
=$\mathcal{R}=\bigg{\{}r:\underset{s_{0}\sim\rho,\atop a_{i}\sim\mathcal{\pi}_{\text{cal}}(\cdot|s_{i})}{{\mathbb{E}}}\sum_{i=0}^{H-1}r(s_{i},a_{i})=0\bigg{\}}.$ (34)
Combining (26) with (29), similar to previous derivations, we have
$J^{\star}(r)$ $=\underset{s_{0}\sim\rho}{{\mathbb{E}}}\left[V^{\star}_{\beta}(s_{0})\right]$
$=\underset{s_{0}\sim\rho,\atop a_{i}\sim\mathcal{\pi}_{\text{cal}}(\cdot|s_{i})}{{\mathbb{E}}}\left[V^{\star}_{\beta}(s_{0})\right]$
$=\underset{s_{0}\sim\rho,\atop a_{i}\sim\mathcal{\pi}_{\text{cal}}(\cdot|s_{i})}{{\mathbb{E}}}\left[\sum_{i=0}^{H-1}r(s_{i},a_{i})-\beta\sum_{i=0}^{H-1}\log\frac{\pi_{r}(a_{i}|s_{i})}{\pi_{\text{ref}}(a_{i}|s_{i})}\right]$
$=-\beta\underset{s_{0}\sim\rho,\atop a_{i}\sim\mathcal{\pi}_{\text{cal}}(\cdot|s_{i})}{{\mathbb{E}}}\left[\sum_{i=0}^{H-1}\log\frac{\pi_{r}(a_{i}|s_{i})}{\pi_{\text{ref}}(a_{i}|s_{i})}\right],$ (35)
where the penultimate line uses (4), and the last line uses Assumption 3.
Consequently, the token-level VPO can be rewritten as
$\pi_{\mathsf{VPO}}$ $=\arg\min_{\pi}\Bigg{\{}-\sum_{(\tau_{+},\tau_{-})\in\mathcal{D}}\log\sigma\left(\beta\sum_{i=0}^{H-1}\log\frac{\pi(a_{i}^{+}|s_{i}^{+})}{\pi_{\text{ref}}(a_{i}^{-}|s_{i}^{-})}-\beta\sum_{i=0}^{H-1}\log\frac{\pi(a_{i}^{-}|s_{i}^{-})}{\pi_{\text{ref}}(a_{i}^{-}|s_{i}^{-})}\right)$
$\qquad\qquad\qquad+\mathsf{sign}\cdot\alpha\beta\underset{s_{0}\sim\rho,\atop a_{i}\sim\mathcal{\pi}_{\text{cal}}(\cdot|s_{i})}{{\mathbb{E}}}\left[\sum_{i=0}^{H-1}\log\frac{\pi(a_{i}|s_{i})}{\pi_{\text{ref}}(a_{i}|s_{i})}\right]\Bigg{\}}.$ (36)
@@ A.2 Proof of Lemma 2
Lemma 5 ((Zhang,, 2023, Theorem 13.2),(Liu et al.,, 2024, Lemma D.1))